OpenStudy (anonymous):

What is the solution to the equation log 3 x - log 3^2 = 2 ? x=18 x=12 x=9/2 x=4/3

OpenStudy (australopithecus):

you Need to apply two rules to simplify this problem Notice that 1. log(x^(y)) = (y)*log(x) 2. log(x) - log(y) = log(x/y) Also remember that 10^log(x) = x Thus by making both sides of the equation exponents to 10 you can eliminate the log and solve for x

OpenStudy (australopithecus):

1. Simplify (combine logarithms) 2. Eliminate log 3. Solve for x If you get stuck post and I will help

OpenStudy (anonymous):

im stuck :/ im not really sure what to do next?

OpenStudy (australopithecus):

what have you done so far?

OpenStudy (australopithecus):

sorry for going something came up that I had to deal with

OpenStudy (australopithecus):

Look at this rule: 2. log(x) - log(y) = log(x/y) Now Look at your problem log(3x) - log(3^2) = 2

OpenStudy (australopithecus):

The way to simplify this is staring you in the face now :)

OpenStudy (australopithecus):

remember x and y are any numbers or variables

OpenStudy (australopithecus):

log(3x) - log(3^2) = 2 = log(3x/3^(2)) = 2 = 10^(log(3x/3^(2))) = 10^(2) = 3x/3^(2) = 10^(2) If you can solve this you can get the answer or conversely you can cut and paste this into https://www.wolframalpha.com/ to get the answer. Remember though by not learning this stuff you are only hurting yourself especially if you are under the age of 25.

OpenStudy (anonymous):

thank you sorry i was gone when u replied :P